### v2

# XAL: EXplainable Active Learning Makes Classifiers Better Low-resource Learners

Abstract
Active learning (AL), which aims to construct an effective training set by iteratively curating the most formative unlabeled data for annotation, has been widely used in low-resource tasks. Most active learning techniques in classification rely on the model’s uncertainty or disagreement to choose unlabeled data, suffering from the problem of over-confidence in superficial patterns and a lack of exploration. Inspired by the cognitive processes in which humans deduce and predict through causal information, we take an initial attempt towards integrating rationales into AL and propose a novel Explainable Active Learning framework (XAL) for low-resource text classification, which aims to encourage classifiers to justify their inferences and delve into unlabeled data for which they cannot provide reasonable explanations. Specifically, besides using a pre-trained bi-directional encoder for classification, we employ a pre-trained uni-directional decoder to generate and score the explanation. We further facilitate the alignment of the model with human reasoning preference through a proposed ranking loss. During the selection of unlabeled data, the predicted uncertainty of the encoder and the explanation score of the decoder complement each other as the final metric to acquire informative data. Extensive experiments on six datasets show that XAL achieves consistent improvement over 9 strong baselines. Analysis indicates that the proposed method can generate corresponding explanations for its predictions.
XAL: EXplainable Active Learning Makes Classifiers Better Low-resource Learners
[…]
1 Introduction
Active learning (AL) is a machine-learning paradigm that efficiently acquires data for annotation from a (typically large) unlabeled data pool and iteratively trains models (Lewis and Catlett, 1994; Margatina et al., 2021). AL frameworks have attracted considerable attention from researchers due to their high realistic values reduce the data annotation costs by concentrating the human labeling effort on the most informative data points, which can be applied in low-resources tasks (Lewis and Catlett, 1994; Settles, 2009; Zhang et al., 2022b).
Most previous AL methods rely on model predictive uncertainty or disagreement for unlabeled data, and the most uncertain data are believed to be the most informative and worthful ones to be annotated (Lewis, 1995; Houlsby et al., 2011; Margatina et al., 2021; Zhang et al., 2022a). However, previous studies have indicated that existing models struggle to accurately quantify predictive uncertainty (Guo et al., 2017; Lakshminarayanan et al., 2017), leading to overconfidence and insufficient exploration, i.e., models tend to choose data instances that are uncertain yet repetitively uninformative (Margatina et al., 2021). This issue arises because training can lead cross-entropy-based classifiers to learn superficial or spurious patterns (Guo et al., 2022, 2023; Srivastava et al., 2020), rather than the causal information between inputs and labels.
Figure 1: Data selection strategy in AL. Previous work selects the unlabeled data mostly relying on the model’s uncertainly (a), but we propose to further leverage the model’s explanation of its prediction (b).
In the context of cognitive science and psychological science, humans make decisions or inferences by exploring causal information (Frye et al., 1996; Joyce, 1999; Rottman and Hastie, 2014). For example, when learning to differentiate animals, humans do not merely rely on statistical features such as colors or feathers. They also consider the creatures’ habits, such as dietary patterns, and kinship, such as the species of the parents, to engage in exploring rationales, thereby determining the species of the organism. Intuitively, explanations of the rational can help the model confirm whether it understands how to make classifications, and explaining the reasons behind the classification also enhances the justification of the inference confidence. It motivates us to encourage classifiers to learn the rationales behind inferences and explore unlabeled data for which the model cannot provide reasonable explanations. In doing so, the model can learn rationales between labels and texts and reduce reliance on superficial patterns, which leads to improved generalization and more effective exploration within AL frameworks. The intuition is illustrated in Figure 1.
Given the above observations, we introduce an Explainable Active Learning Framework (XAL) for text classification tasks. This framework consists of two main components: the training process and the data selection process. Primarily, we adopt a pre-trained bi-directional encoder for classification and a pre-trained uni-directional decoder to generate and score explanations that serve as expressions of rationales in human language. In the training phase, we use the classification labels and explanations to optimize the model parameters. Besides, to further enhance the decoder’s ability to score explanations, we design a ranking loss that optimizes the model to differentiate between reasonable and unreasonable explanations. To implement this ranking loss, we generate a variety of explanations (both reasonable and unreasonable) for labeled data by querying ChatGPT with different prompts, thereby eliminating the need for additional human annotation effort. Subsequently, during the data selection phase, we amalgamate the predictive uncertainty of the encoder and the explanation score of the decoder to rank unlabeled data. The most informative data are then annotated and incorporated into further training.
We conduct experiments on various text classification tasks involving different level of difficulty in understanding rationals. Experimental results manifest that XAL can achieve substantial improvement in all tasks. Ablation studies demonstrate the effectiveness of each component, and human evaluation shows that the model trained in XAL works well in explaining its prediction. XAL also demonstrates superior performance with only 500 instances when compared to in-context learning by ChatGPT, underscoring the effectiveness of our model at a minimal cost. To our knowledge, we are the first to incorporate the model’s explanation (explanation score) to improve the effectiveness of data selection in AL process. We will release the code and data to facilitate further research.
2 Related Work
Figure 2: Our proposed XAL framework, which can be divided into two main parts – the training process (red arrows) and the data selection process (blue arrows). The training process aims to train the encoder-decoder model to learn classification and explanation generation. The data selection process aims to select unlabeled data using predictive entropy and explanation scores.
Active Learning is widely studied in the natural language processing area, ranging from text classification (Roy and McCallum, 2001; Zhang et al., 2017; Maekawa et al., 2022), and sequence labeling (Settles and Craven, 2008) to text generation (Zhao et al., 2020). Previous methods can be roughly divided into informativeness-based selection strategies, representativeness-based selection strategies, and hybrid selection strategies (Zhang et al., 2022a). The most mainstream methods, i.e., informativeness-based methods, are mostly characterized using model uncertainty, disagreement, or performance prediction, which suffers from over-confidence and a lack of exploration (Guo et al., 2017; Margatina et al., 2021). On the other hand, the representativeness-based methods rely on model inputs such as the representations of texts, which tends to select simple data samples and results in unsatisfactory performance (Roy and McCallum, 2001; Margatina et al., 2021).
Large Language Model. Recently, LLMs of generative schema have shown excellent performance in various NLP tasks. However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021). Considering the problems, LLMs are also applied to improve the generalization of smaller models for specific tasks. Some studies distill the knowledge of large language models to a smaller one Hsieh et al. (2023) and some also use LLMs to augment data to achieve stronger text classification performance Ye et al. (2022); Yu et al. (2023). LLMs has also demonstrated a strong capability in generating high-quality reasoning steps Hsieh et al. (2023); Wei et al. (2022); Kojima et al. (2022). In this study we aim to distill the reasoning ability of LLMs to smaller models, and encourage the models to distinguish the reasonablity of explanations to identify informative data in AL scenario.
Explanation Information, as external knowledge, has been proven useful for a wide range of tasks in natural language processing (Hase and Bansal, 2022). Hase et al. (2020) used explanations as additional information and directly fed them into models. Narang et al. (2020) and Shen et al. (2023) took the explanations as outputs and trained NLP models to generate them. How to leverage explanations is still an open problem (Hase and Bansal, 2022). In the active learning schema, some studies also attempt to leverage the explanations (Liang et al., 2020; Wang et al., 2021), but they mainly focus on promoting the generalization abilities of models trained on low-resource data. These AL studies are also hard to implement in text classification tasks and unlike these studies, we explore how to leverage explanations to identify informative unlabeled data for annotation.
3 Method
3.1 Overview
Task Formulation
We mainly consider a $C$ class text classification task defined on a compact set $\mathcal{X}$ and a label space $\mathcal{Y}=\{1,...,C\}$. The data points are sampled i.i.d over the space $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ as $\{\textbf{x}_{i},y_{i}\}\sim p_{z}$, which can be divided into two sets – the labeled set $D_{l}$ and the unlabeled set $D_{u}$. At the beginning of an active learning algorithm, only a small number of data points are randomly selected into the labeled set $D_{l}$ and we have only access to data points in $D_{l}$ for training the classification model. Then $L$ data from $D_{u}$ are selected for annotation and added to $D_{l}$ (removed from $D_{u}$ simultaneously) in $\mathcal{M}$ multiple rounds.
Model Architecture
Following previous work (Devlin et al., 2018), we adopt a pre-trained bi-directional encoder as the backbone for classification. In addition to the encoder, a corresponding uni-directional decoder is applied to generate and score the explanation for the label prediction. During training, we construct $k$ different explanations $\textbf{e}_{r}$, i.e., $\{\textbf{e}_{r}\}_{i},r=0,...,k-1,$ for each example $\{\textbf{x}_{i},y_{i}\}$, where $\textbf{e}_{0}$ is the reasonable explanation and $\{\textbf{e}_{r>0}\}$ are $k-1$ unreasonable explanations. We leave the construction process of explanations in Section 3.4 for further descriptions. Before that, we will first present the model training and data selection in Section 3.2 and Section 3.3 respectively. The framework of XAL is shown in Figure 2 and the workflow can be found in Algorithm 1.
3.2 Training
For each text input x (we omit all the subscripts of $i$ for simplicity in this subsection), we first prepend it with a special token $[CLS]$ and then obtain the contextual representation by feeding it into the encoder. The contextual representation of the $j$th token is calculated as:
$\textbf{h}_{j}=Encoder([CLS]+\textbf{x})[j].$ (1)
The representation for [CLS], i.e., $\textbf{h}_{0}$ is taken as the sentence representation and fed into the classification layer, which is composed of a linear layer and a softmax function. The probability distribution on label space $\mathcal{Y}$ can be formulated as:
$P(y|\textbf{x})=Softmax(Linear(\textbf{h}_{0})).$ (2)
The cross-entropy loss is adopted to optimize the encoder parameters:
$\mathcal{L}_{cls}=-\sum P(y|\textbf{x})\ log\ P({y}|\textbf{x}).$ (3)
On the decoder side, the model is trained with teacher forcing to generate the golden explanation $\textbf{e}_{0}$. The generation loss is calculated as:
$\mathcal{L}_{gen}=-\sum_{t}log\ P(\textbf{e}_{0,t}|\textbf{h},\textbf{e}_{0,<t}).$ (4)
To make the decoder a good scorer to rank the reasonable and unreasonable explanations, we additionally adopt a ranking loss to optimize the decoder. In particular, the model is trained to rank between reasonable and unreasonable explanations. The ranking loss can be formulated as:
$\mathcal{L}_{rank}=\sum_{r>0}max(0,p_{r}-p_{0}),$ (5)
where $p_{r}$ is the explanation score for $\textbf{e}_{r}$, calculated as the length-normalized conditional log probability:
$p_{r}=\frac{\sum_{t}log{P(\textbf{e}_{r,t}|\textbf{x},\textbf{e}_{r,<t})}}{||\textbf{e}_{r}||}.$ (6)
The hyper-parameters are adopted to balance the weights of each loss, and the overall loss is formalized as follows:
$\mathcal{L}=\mathcal{L}_{cls}+\lambda_{1}\mathcal{L}_{gen}+\lambda_{2}\mathcal{L}_{rank}.$ (7)
3.3 Data Selection in AL
After training the model in each iteration, we can obtain an intermediate model $\pi$. To select the informative data in the unlabeled set $\mathcal{D}_{u}$, we adopt a combination of the predictive entropy and explanation score. Specifically, for each raw data $\textbf{x}_{i}\in\mathcal{D}_{u}$, we first generate the explanation $\textbf{e}_{i}$ by selecting the top-$1$ output in the beam search. Then, we calculate the explanation score $p_{i}$ as Eq. 6 and the predictive entropy $c_{i}$ as Eq. 3. The final score $s_{i}$ for example $\textbf{x}_{i}$ is calculated as the weighted sum of the normalized explanation score and predictive entropy:
$s_{i}=\frac{\lambda}{1+\lambda}\frac{e^{-p_{i}}}{\sum_{i}e^{-p_{i}}}+\frac{1}{1+\lambda}\frac{e^{c_{i}}}{\sum_{i}e^{c_{i}}}$ (8)
where the $\lambda$ is the hyper-parameter to balance the explanation score and the predictive entropy. With the final score for each example, we rank the whole unlabeled instances and select the top $L$ instances for annotation.
3.4 Generation of Golden Explanations
Previous work has shown that LLMs are good at reasoning (Bang et al., 2023; Rajasekharan et al., 2023). Inspired by these studies, we take the LLMs, such as ChatGPT and GPT4, as the teacher models, and query them to generate explanations for each selected labeled data, eliminating the annotation cost of human labor. In particular, we design slightly different prompt templates for different tasks, and the prompt for each task is shown in Appendix A. Taking stance detection as an example, its prompt template is designed as ‘The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words’, where the Target is the corresponding stance target, and the Label is the classification label. The final query to the teacher model is the concatenation of the text and the prompt. We construct a reasonable explanation by feeding the golden label into the query and generate several unreasonable explanations by feeding wrong labels. Figure 7 shows an example that we generate explanations by querying ChatGPT, where we can observe that ChatGPT could provide different explanations according to the given label.
4 Experiments
Figure 3: Results given the data selection budget 500 instances in six text classification tasks, where 100 instances are selected for annotation in each iteration. Here we plot the specific values of XAL and the second significant performance when using 500 instances, and the detailed performance values can be found in Appendix E.1.
4.1 Tasks and Dataset
We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text; (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff; (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target; (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment; (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic. We adopt the datasets RTE Bentivogli et al. (2009), MRPC Dolan et al. (2004), MAMS Jiang et al. (2019), COVID19 Glandt et al. (2021), DEBA Pougué-Biyong et al. (2021) and CLEF Kanoulas et al. (2017) for each task, respectively. Appendix A demonstrates the details of six datasets with examples.  [^1] Without losing generality, we randomly split the training set in RTE, and MRPC into train/dev set with proportion 9:1. In DEBA, we adopt the topic of climate change for experiments.
4.2 Baselines
To demonstrate the effectiveness of our proposed method, we compare XAL with the following nine AL baselines: (1) Random uniformly selects unlabeled data for annotation; (2) Max-Entropy (ME) (Lewis, 1995; Schohn and Cohn, 2000) calculates the predictive entropy in the current model and selects data with max entropy ; (3) Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011) exploits the uncertainty of unlabeled data by applying different dropouts at test time; (4) Breaking Ties (BK) (Scheffer et al., 2001) selects instances with the minimum margin between the top two most likely probabilities ; (5) Least Confidence (LC) (Culotta and McCallum, 2005) adopts instances whose most likely label has the least predictive confidence; (6) Coreset (Sener and Savarese, 2018; Chai et al., 2023) treats the representations in $D_{u}$ as cluster centers, and selects the unlabeled data with the most significant distance from its nearest centers; (7) Batch Active learning by Diverse Gradient Embeddings (BADGE) Ash et al. (2019) measures uncertainty as the gradient magnitude and collects examples where these gradients span a diverse set of directions; (8) Bayesian Estimate of Mean Proper Scores (BEMPS) Tan et al. (2021) encourages diversity in the vector of expected changes in scores for unlabelled data; (9) Contrastive Active Learning (CAL) (Margatina et al., 2021) selects instances with the maximum mean Kullback-Leibler (KL) divergence between its $m$ nearest neighbors.
5 Results and Discussion
Figure 4: Experimental results demonstrate how much data, when selected using AL methods, is required for the models to achieve 90% of the performance of those trained on the complete training datasets. In each iteration, we annotate 50 instances. The performance of models trained on the whole training sets is, (a) RTE – 83.11%, (b) MRPC – 84.74%, (c) COVID19 – 75.45%, and (d) DEBA – 65.71%. The green triangles refer to the average values of the experiments on three different initial sets $D_{l}$ and three different random seeds. The circles refer to outliers. Detailed results can be seen in Appendix E.2.
5.1 Main Results
We mainly consider two different settings: (1) Given the data selection budget, we observe the trend of changes in model performance; (2) Given the performance upper bound, we observe the number of required instances that the model needs to achieve 90% of the upper-bound performance. We utilize FLAN-T5-large Chung et al. (2022) as our backbone network 222 We implement baselines using FLAN-T5-large as well., ChatGPT is adopted to generate the explanations. The implemented details can be found in Appendix D and the detailed values of the result can be found in Appendix E.1.
Given Data Selection Budget
Following previous work (Zhang et al., 2017; Schröder et al., 2022), we set the data selection budget as 500 instances and select 100 instances for annotation in each iteration. The results are presented in Figure 3. We can observe that the proposed XAL model consistently outperforms other active learning methods. For instance, in MAMS, our model attains a macro-F1 score of 74.04% at the end, which is 2.21% higher than the second-best result (LC at 71.83%). Similarly, in DEBA, XAL surpasses the second-best result (BADGE at 58.13%) by 1.78%. These results demonstrate the effectiveness of our XAL framework in addressing text classification tasks.
In COVID19, while the model does not significantly outperform the baselines at the beginning (possibly due to the relatively high complexity of the task), it still exhibits stronger performance with a data count of 300-500, which underscores the effectiveness of the data selection in XAL. In CLEF, we notice that the performance of baseline models is notably unstable due to the significant imbalance in label distribution (the ratio between relevant and irrelevant is approximately 1:21). However, our XAL model achieves superior performance and more consistent improvements over baselines during the data selection process, which validates the effectiveness of XAL, even in challenging scenarios of imbalanced data.
Given Performance Upper Bound
It’s also valuable to evaluate the amount of data required for models to achieve comparable performance with those trained on the entire training dataset. Specifically, we begin with an initial labeled set of 100 instances and select a certain number of instances to annotate in each selection iteration  [^3] To balance the training efficiency and the performance gap, we set the selection number as 50., and cease the AL process once the model performance reaches 90% of the upper-bound performance. Experimental results are depicted in Figure 4.444For ease of presentation and training efficiency, we only report results on four tasks. As observed, XAL requires the least amount of data to reach the performance goal. For instance, in the task of DEBA, XAL necessitates an average of 461.11 data points, which is 55.56 less than the second lowest value (BK–516.67). To conclude, XAL models only require 6%, 3%, 16%, and 10% of the data from RTE, MRPC, COVID19, and DEBA tasks respectively to achieve 90% performance of models that are trained on the entire datasets, which significantly reduces the annotation cost. These results show that the proposed XAL is very cost-efficient in selecting informative unlabeled data.
Figure 5: Results of ablation study in the six text classification tasks. We select 100 instances in each iteration and conduct 4 iterations (the same with Section 5.1). The results are measured using macro-F1 scores and they are the average values on three different initial sets $D_{l}$ and three different random seeds.
5.2 Ablation Study
We conduct an ablation study to investigate the impact of each module in our model. The results are displayed in Figure 5. Firstly, we conduct a comparison among ME, ME-Exp, and XAL, where ME-Exp has the same model structure as XAL but it selects the unlabeled data with the predicted classification entropy. We observe that ME-Exp can achieve superior performance on most datasets compared to ME, which demonstrates the effectiveness of using explanations. However, XAL further achieves noticeably better performance over ME-Exp, indicating that the improvement in XAL comes not only from the introduction of explanations but also from the data selection strategy (with explanation scores). Next, we compare XAL with a version that removes the ranking loss (w/o Rank in Figure 5). XAL also achieves better performance on most datasets and with different numbers of labeled data, indicating that the ranking loss can enhance the effectiveness of data selection in the AL process. Furthermore, the performance of selecting data solely using the explanation score but without using predictive entropy is also illustrated in Figure 5 (w/o ME). We observe that removing ME leads to significant performance drops on most datasets, implying that the predictive entropy and explanation score can complement each other.
100 200 300 400 500
ChatGPT 60.79 63.47 68.51 71.54 73.24
ALPACA-7B 59.52 61.75 67.77 71.12 72.24
GPT4 59.67 64.28 69.51 72.96 74.63
Table 1: Model performance on MAMS using different explanation generations. We compare the performance in a certain initial set and random seed.
MRPC COVID19 DEBA CLEF
Zero-shot 72.46 66.67 48.96 34.21
Few-shot 78.32 67.73 54.69 42.44
Silver 74.32 53.66 47.12 36.09
XAL (500) 83.91 67.16 59.91 66.94
Table 2: Performance of ChatGPT and performance of models trained on ‘silver’ data.
To further evaluate how the ranking loss works in XAL, we also compare the model’s capability to rank explanations between XAL and its counterpart without ranking loss. Experimental results show that XAL achieves superior performance. For instance, the ranking accuracy in RTE and MRPC for XAL are 73.93% and 78.62%, which are 5.36% and 4.30% higher than those without ranking loss, respectively (detailed results are shown in Appendix E.4). These results suggest that the ranking loss can enhance the model’s ability to score the explanations. It is evident that XAL consistently outperforms these alternatives in most time, while there are some fluctuations across different scenarios.
5.3 Explanation Generation
We also carry out experiments to analyze how the generation of explanations impacts model performance. Specifically, we replace ChatGPT with ALPACA-7B (Taori et al., 2023), and GPT4  [^5] https://openai.com/gpt-4 to generate explanations on the MAMS dataset. The results are presented in Table 1. We also observe that the ALPACA-7B can also provide useful explanations to some extent and enhance the model performance compared with ME. This suggests that LLMs, when used as an assistant in XAL, can provide consistent explanation generation and enhance model performance. The results also indicate that the model performance can be affected by the explanation generation model and it is applicable to use open-source LLMs. The results of human annotation are also discussed in Appendix F.
5.4 Comparison with ChatGPT
We assess ChatGPT on these datasets in both zero-shot and few-shot scenarios, and the outcomes are presented in Table 2. The models, when subjected to supervised fine-tuning with 500 labeled data points in XAL, exhibit either notably improved or comparable performance to ChatGPT across these datasets. This suggests that our model can achieve satisfactory results with minimal cost.
Since XAL queries ChatGPT for generating explanations with extra cost (API calls) to implement AL, we also analyze if we could obtain a better model through querying ChatGPT for more labeled data, i.e. ‘silver data’ (the labels are probably incorrect). In detail, we use the actively selected data with golden labels, and randomly selected data with ‘silver’ labels from ChatGPT to train the model in each active iteration. For this process, ChatGPT is employed to annotate a random selection of data from $D_{u}$, which is three times the size of $D_{l}$—mirroring the frequency of queries made for explanations in XAL. We report the final model performance with 500 golden labels and 1500 silver labels in Table 2. We observe that despite the increase in silver data obtained from ChatGPT, our model can still perform more significantly. This outcome is primarily attributed to the uncertain accuracy of ChatGPT’s annotations. The lack of reliability in these pseudo labels suggests that merely increasing the quantity of labeled data, without ensuring its quality, may not be an effective strategy for improving model performance.
5.5 Human Evaluation on Interpretability
We evaluate our model’s ability to explain its prediction by examining the consistency between the generated explanation and the classification label. Specifically, we randomly select 50 test instances and use the model trained on 500 instances (see Section 5.1) to generate the labels and explanations. Then we ask humans to infer the classification labels based solely on the generated explanations. The consistency is measured by whether the human-inferred label equals the label predicted by the model. We report the consistency rate across all the test sets: MRPC-94%, RTE-94%, COVID19-96%, DEBA-94%, MAMS-94%, CLEF-100%. We find that the consistency rates on all six tasks exceed 94%, which demonstrates that XAL explains its classification prediction very well. Case studies for the generated explanations and the predicted labels are presented in Appendix I.
Figure 6: t-SNE visualizations of contextual representations. To facilitate identification, we outline the areas of labeled data with dashed squares. The colors, i.e. red, blue, and green, refer to the different labels.
5.6 Representation Visualization
To understand the potential of XAL in exploring informative unlabeled data, we use t-SNE (van der Maaten and Hinton, 2008) to “visualize” the data selection procedure of ME and XAL on the task DEBA. Specifically, with the intermediate model in Section 5.1 (trained with 200 labeled instances), 100 instances from the unlabeled set $D_{u}$ are then selected for annotation. Then, we feed all the labeled and unlabeled instances into the model and get their sentence representations ($\textbf{h}_{0}$ in Eq. 1). Finally, we apply the t-SNE toolkit to map these representations into a two-dimensional embedding space, which is shown in Figure 6. We can observe that the data are obviously partioned into different clusters, which shows the overconfidence problem in ME, and the unlabeled data selected by ME is only distributed around the decision boundary, showing that the model can only select the high-uncertainty data it believes. However, the proposed XAL can select more diverse data, some of which are wrongly classified by the current model. These results demonstrate that the data selection strategy in XAL can identify more informative data and mitigate the problem of overconfidence to some extent. More visualizations are shown in Appendix J.
6 Conclusion
In this paper, we proposed a novel Explainable Active Learning (XAL) framework for text classification. Experiments demonstrated that XAL achieves substantial improvements compared with previous AL methods. Further analysis indicated that the proposed method can generate corresponding explanations for its predictions.
7 Limitations
XAL takes an initial attempt towards integrating rationales information into active learning. While acknowledging that this approach may necessitate additional computational resources, this augmentation empowers the trained classifier to be both more explainable and more generalized, as the model can generate explanations for its predictions and obtain enhanced performance. Our model, which incorporates a decoder module to obtain the generation score, necessitates more time for data selection, which is detailed in Appendix G, but during the inference, since we use an encoder-decoder models for training, we can directly use the encoder for inference if the explanation generation is not in need. In our experiments, we evaluated our model’s effectiveness across six classification tasks in a low-resource setting, but XAL can be used for other tasks with more label classes and industrial downstream applications.
8 Ethical Statement
We honor the ACL Code of Ethics. No private data or non-public information was used in this work. All annotators have received labor fees corresponding to the amount of their annotated instances.
Appendix A Tasks and Corresponding Prompts
We first introduce the dataset and its statistics in Table 3. We show the tasks and examples for experiments in Table 4, including natural language inference, paraphrase detection, category sentiment classification, stance detection, (dis)agreement detection, and relevance classification. Then we show how we obtain different explanations in Figure 7 with an example of COVID19. We also show the prompts we used for explanation generation through querying ChatGPT (Table 5).
Figure 7: The process to generate diverse explanations from LLMs. We can obtain reasonable and unreasonable explanations by querying ChatGPT with correct and incorrect labels, respectively.
Task
Dataset
# Labels Train Dev Test
Natural Language Inference
RTE (Bentivogli et al., 2009)
2 2,240 250 278
Paraphrase Detection
MRPC (Dolan et al., 2004)
2 3,667 409 1,726
Stance Detection
COVID19 (Glandt et al., 2021)
3 4,533 800 800
Category Sentiment Classification
MAMS (Jiang et al., 2019)
3 7,090 888 901
(Dis)agreement Detection
DEBA (Pougué-Biyong et al., 2021)
3 4,617 578 580
Relevance Classification
CLEF (Kanoulas et al., 2017)
2 7,847 981 982
Table 3: All the six text classification tasks used in our experiments. The extent of difficulty is roughly in an increasing tendency. RTE and MRPC are fundamental natural language tasks and are included in the widely used benchmark GLUE. MAMS, COVID19, and DEBA require the model to understand the text and give suitable inferences towards a specific target or text, and CLEF further provides a difficult dataset with imbalanced label distribution.
Task
Text
Natural Language Inference
Sentence 1: Danny Kennedy, Greenpeace campaigns director, said: "The burden of proof in the Scott Parkin expulsion case lies morally with the Commonwealth, to prove that he is a danger. When the Government brought in anti-terror legislation, they promised the public that these laws would only be used to confront a real and present risk of a terrorist attack, not a sweep-all approach against citizens. Peace is not terrorism. Peace is not a threat to national security. No democratic government should expel a foreign citizen because [it] opposes his political opinions."
Sentence 2: Greenpeace director said that peace is terrorism.
Label: Not Entailment.
Paraphrase Detection
Sentence 1: Last week the power station’s US owners, AES Corp, walked away from the plant after banks and bondholders refused to accept its financial restructuring offer .",
Sentence 2: "The news comes after Drax’s American owner, AES Corp. AES.N, last week walked away from the plant after banks and bondholders refused to accept its restructuring offer.
Label: Paraphrase/Semantic Equivalent.
Category Sentiment Classification
Text: I left feeling unsatisfied, except for having a nice chance to people watch in the cozy atmosphere with my over-priced pasta bolognese.
Target: Ambience
Label: Positive
Stance Detection
Text: Michigan is fining individuals 500\$ for not wearing a mask in public. How do y’all feel about this? Curious because I am torn about being so forceful but agree that people should wear masks. #MaskOn.
Target: Face Mask
Label: Favor
(Dis)agreement Detection
Text 1: True, but with lower power usage, you have less heat to dissipate, meaning you can overclock it even more.
Text 2: AMD creates a chip that saves energy by over 31 times. Someone show this to r/PCMasterRace cause we need to switch to AMD.
Label: Agree.
Relevance Classification
Document: 99mtechnetium penicillamine: a renal cortical scanning agent. 99mTechnetium penicillamine, a renal cortical imaging agent, can be used to provide a rapid, safe, and non-invasive assessment of renal morphology and the renal vascular supply. Since this agent is not excreted significantly during the imaging procedure cortical scans of high quality can be obtained without image deterioration owing to a superimposed collecting system. These scans, which are clearly superior in anatomical detail to earlier scans using 131I hippuran, can be obtained along with the 131I hippuran renogram when the patient comes to the nuclear medicine department. Herein we demonstrate the anatomical detail it is now possible to achieve by presenting the cortical renal scans and accompanying radiograms from 5 patients with different renal pathology.
Topic: Procalcitonin, C-reactive protein, and erythrocyte sedimentation rate for the diagnosis of acute pyelonephritis in children.
Label: Not Relevant.
Table 4: Tasks and examples for experiments.
Task
Prompts
Label set
Natural Language Inference
Sentence 1: {Text 1}. Sentence 2: {Text 2}. Sentence 1 can {Label} sentence 2, explain the reason within 50 words.
{Entail, Not Entail}
Paraphrase Detection
Sentence 1: {Text 1}. Sentence 2: {Text 2}. The relation between the above two sentences is {Label}, explain the reason within 50 words.
{Paraphrase Equivalent, Not Paraphrase Equivalent}
Category Sentiment Classification
{Text}.
{Positive, Negative, Neutral}
Stance Detection
{Text}. The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words.
{Favor, Neutral, Against}
(Dis)agreement Detection
Sentence 1: {Text 1}. Sentence 2: {Text 2}. The stance of sentence 2 is {Label} to sentence 1, explain the reason within 50 words.
{Agree, Neutral, Disagree}
Relevance Classification
{Document}. The scientific document is {Label}’ to the research topic {Topic}, explain the reason within 100 words.
{Relevant, Not Relevant}
Table 5: The prompts we adopted for the tasks in our experiments.
Appendix B Explanation Examples
Using the prompts in Appendix A, we show some examples of the obtained explanations in Table 6 by querying ChatGPT.
Input
Label
Explanation
RTE
Sentence 1: The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes.
Sentence 2: Nobel Peace Prize candidates have been chosen.
Not Entailment
Sentence 1 states that the Nobel prizes are the main forms of recognition for various fields, including peace. However, sentence 2 only mentions that candidates have been chosen for the Nobel Peace Prize, without specifying if they have actually received the prize. Therefore, sentence 2 does not necessarily follow from sentence 1.
Entailment
Sentence 1 can entail sentence 2 because sentence 1 mentions that one of the main forms of recognition for those working for peace is the Nobel prizes. Therefore, it is reasonable to assume that candidates for the Nobel Peace Prize have been chosen.
MRPC
Sentence 1: Maddox, 87, cracked two ribs when he fell about 10 days ago at an assisted living home where he was recovering from intestinal surgery, Virginia Carnes said.
Sentence 2: Maddox, who had battled cancer since 1983, cracked two ribs earlier this month when he fell at an assisted living home where he was recovering from surgery.
Equivalent
The two sentences provide similar information but are expressed differently. Sentence 1 mentions that Maddox fell about 10 days ago, while Sentence 2 specifies that the fall occurred earlier this month. Despite the slight difference in wording, both sentences convey the same meaning.
Not Equivalent
The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information about the timing and reason for Maddox’s rib injury. In the first sentence, it mentions that Maddox fell about 10 days ago while recovering from intestinal surgery, whereas the second sentence states that he fell earlier this month while recovering from surgery without specifying the exact time frame.
COVID19
Tweet: Placed2nd @AngelaBelcamino Omg people say I have to wear a mask and I’m crying because I have to wear a mask to save lives boo hoo!! MAGAts are stupid!!!
Favor
The stance of this tweet towards face masks is favorable. The reason for this is that the tweet expresses support for wearing masks to save lives and criticizes those who oppose wearing masks, referring to them as ""MAGAts"" (a term often used to describe supporters of former President Donald Trump)."
Against
The reason behind the stance against face masks in this tweet seems to be frustration and ridicule towards those who oppose wearing masks. The tweet suggests that the person finds the idea of wearing masks to save lives unnecessary and mocks those who support it, particularly referring to them as ""MAGAts,"" a derogatory term for supporters of former President Trump.
Neutral
The tweet expresses a neutral stance towards face masks by mocking the idea of crying over having to wear one. While it criticizes MAGA supporters, it does not explicitly state a negative or positive opinion about face masks themselves."
Table 6: Diverse explanations generated from ChatGPT. The golden label is marked in bold.
Appendix C Algorithm
We show the detailed algorithm of XAL in Algorithm 1.
Algorithm 1 Explainable Active Learning Algorithm
1:  Initialization: dataset $D_{u}$, iteration steps $\mathcal{M}$, selective number $L$, training epoch $\mathcal{T}$.
2:  Randomly select $L$ data from $D_{u}$, denoted as $D_{s}$ and remove them in $D_{u}$.
3:  Annotate the data $\textbf{x}_{i}\in D_{s}$ for $y_{i}^{c}$ with human annotators.
4:  Query ChatGPT for diverse explanations $y_{i}^{g_{r}}$ for the data $\{\textbf{x}_{i},y_{i}^{c}\}\in D_{s}$.
5:  Add $\{\textbf{x}_{i},y_{i}^{c},y_{i}^{g_{r}}\}\in D_{s}$ to $D_{l}$, and empty the set $D_{s}$.
6:  $m=1$.
7:  repeat
8:     $m\leftarrow m+1$
9:     Initialize an explainable classifier $\pi$ and $t=0$.
10:     repeat
11:        $t\leftarrow t+1$
12:        Calculate optimization loss using data $\{\textbf{x}_{i},y_{i}^{c},y_{i}^{g_{r}}\}\in D_{l}$.
13:        Optimize the explainable classifier $\pi$.
14:     until $t>\mathcal{T}$
15:     Calculate the predictive entropy $\textbf{p}_{i}$ and explanation scores $\textbf{c}_{i}$ of data $\textbf{x}_{i}\in D_{u}$ uing Eq. 6.
16:     Calculate the rank score using Eq. 8.
17:     Select $L$ data with the largest score from $D_{u}$ to $D_{s}$.
18:     Annotate the data in $D_{u}$ following the steps 3-5.
19:  until $m>\mathcal{M}$
19:  Explainable classifier $\pi$.
Appendix D Implementation Details
In our experiments, we directly utilize a pre-trained encoder-decoder language model for its strong ability in text understanding and generation. Specifically, we adopt the officially released pre-trained FLAN-T5-Large model (Chung et al., 2022) from Huggingface  [^6] https://huggingface.co/. All models in our experiments are trained on a single GPU (Tesla V100) using the Adam optimizer (Kingma and Ba, 2014). We set the learning rate at 1e-4, with a linear scheduler. The batch size is consistently set to 1 across all tasks. The models are trained for 10 epochs in each iteration. Hyper-parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda$ are set to 0.1, 0.01, and 0.5, respectively, based on preliminary experiments. Note that here we don not specially tune the hyperparameters using grid search, but select the parameters considering their magnitude and keep the same across different tasks, and we further implement sensitivity analysis in Appendix H. The performance for all tasks is evaluated based on macro-averaged F1. The reported results are the average of three initial sets $D_{l}$ and three random seeds (the average of 9 experimental results overall).
Appendix E Detailed Results
E.1 Main Results
The details of the main results are shown in Table 7.
Random ME BALD BK LC Coreset CAL BADGE BMEPS XAL
RTE
100 67.34 67.34 67.34 67.34 67.34 67.34 67.34 67.34 67.34 73.40
200 70.64 73.37 68.06 72.80 72.91 71.12 72.22 72.76 70.36 78.22
300 72.16 76.15 71.09 76.85 75.60 74.90 74.81 78.64 75.53 80.51
400 74.71 78.35 73.11 76.90 78.15 75.16 76.15 80.37 78.29 81.50
500 75.54 77.69 74.30 79.44 79.22 77.69 75.42 81.52 80.66 83.02
MRPC
100 69.80 69.80 69.80 69.80 69.80 69.80 69.80 69.80 69.80 75.31
200 75.44 76.26 73.95 76.35 77.10 76.54 76.22 74.78 76.21 80.73
300 78.12 80.14 76.07 80.23 79.87 79.39 78.52 79.21 80.02 81.31
400 80.28 80.74 77.64 81.95 81.21 79.85 79.76 80.67 81.91 82.76
500 80.63 80.90 79.90 82.33 81.53 80.06 80.60 82.11 82.44 83.91
MAMS
100 56.73 56.73 56.73 56.73 56.73 56.73 56.73 56.73 56.73 59.32
200 61.77 63.01 58.75 62.34 62.83 62.59 61.89 64.57 59.64 66.19
300 66.38 65.90 62.68 66.92 66.72 64.83 65.96 67.18 64.48 69.16
400 67.88 69.44 64.33 69.67 69.74 68.93 67.54 70.26 69.88 71.74
500 70.05 71.23 66.69 71.78 71.83 69.50 69.59 71.35 70.51 74.04
COVID19
100 52.29 52.29 52.29 52.29 52.29 52.29 52.29 52.29 52.29 52.24
200 57.19 54.84 53.19 57.22 55.67 56.18 55.67 55.14 55.62 57.57
300 57.95 59.80 54.74 60.10 58.45 58.18 58.45 59.13 58.67 60.48
400 59.85 61.73 55.98 62.30 61.38 60.62 61.38 61.01 59.63 63.63
500 61.78 64.30 56.01 64.36 64.48 61.45 64.48 63.25 62.88 67.16
DEBA
100 42.09 42.09 42.09 42.09 42.09 42.09 42.09 42.09 42.09 46.21
200 50.60 48.74 46.81 50.65 49.73 49.18 49.26 47.35 49.24 53.16
300 53.93 52.43 51.54 54.87 54.57 53.97 53.43 54.21 55.22 57.35
400 57.03 56.58 53.18 57.02 57.15 56.37 55.06 56.75 57.12 58.03
500 57.45 57.25 55.66 57.78 57.64 56.95 55.82 57.88 58.13 59.91
CLEF
100 57.72 57.72 57.72 57.72 57.72 57.72 57.72 57.72 57.72 60.02
200 57.95 59.50 57.38 59.82 58.44 59.62 60.67 58.42 58.14 60.13
300 57.37 60.23 58.80 60.66 61.72 59.60 58.44 59.88 60.12 61.97
400 60.04 63.52 59.14 63.48 63.81 61.46 59.04 61.34 61.22 63.97
500 61.04 62.57 59.03 64.66 65.16 59.84 61.53 64.12 63.27 66.94
Table 7: Main results in the six text classification tasks. We select 100 instances in each iteration and conduct 4 iterations. The results are measured using macro-F1 scores and they are the average values on three different initial sets $D_{l}$ and three different random seeds.
E.2 Given upper bound
We show the average number of data required for the model to achieve 90% performance of those trained on all the training data (Table 8).
Random ME BK Coreset CAL XAL
RTE 388.89 283.33 255.56 344.44 377.78 133.33
MRPC 227.78 222.22 211.11 216.67 205.56 122.22
COVID19 1305.56 877.78 888.89 1172.22 1088.89 761.11
DEBA 605.56 594.44 516.67 594.44 566.67 461.11
Table 8: The detailed experimental results about how much data queried by AL methods can the model achieve 90% performance of the models trained on the whole training data. In each iteration, we select 50 data. The model performances trained on the whole training sets are, (a) RTE – 83.11%, (b) MRPC – 84.74%, (c) COVID19 – 75.45%, and (d) DEBA – 65.71%. The green triangles refer to the average values of the nine-times experiments.
E.3 Ablation Study
The detailed results of the ablation study are shown in Table 9.
ME ME-Exp w/o Rank w/o ME XAL
RTE
100 67.34 72.09 72.77 71.96 73.40
200 73.37 76.68 76.17 75.10 78.22
300 76.15 80.23 80.22 75.77 80.51
400 78.35 80.05 80.42 78.46 81.50
500 77.69 80.08 82.01 78.36 83.02
MRPC
100 69.80 71.58 75.18 75.64 75.31
200 76.26 77.32 80.37 80.53 80.73
300 80.14 80.93 81.50 82.19 81.31
400 80.74 82.72 82.40 81.61 82.76
500 80.90 82.09 83.02 82.42 83.91
MAMS
100 56.73 59.77 60.69 60.73 59.32
200 63.01 64.57 65.90 65.11 66.19
300 65.90 69.32 69.79 69.32 69.16
400 69.44 71.71 71.38 70.83 71.74
500 71.23 72.97 72.79 71.71 74.04
COVID19
100 52.29 50.83 50.94 51.19 52.24
200 54.84 54.13 55.56 55.25 57.57
300 59.80 58.48 58.51 59.48 60.48
400 61.73 62.63 62.66 61.53 63.63
500 64.30 63.88 64.29 62.90 67.16
DEBA
100 42.11 45.06 47.16 45.48 46.21
200 48.74 51.86 52.26 53.11 53.16
300 52.43 55.74 57.15 56.59 57.35
400 56.58 57.51 57.65 57.35 58.03
500 57.25 58.08 59.09 59.11 59.91
CLEF
100 57.72 57.74 57.91 57.99 60.02
200 59.50 59.24 57.80 58.63 60.13
300 60.23 63.00 61.58 60.58 61.97
400 63.52 63.78 63.40 61.87 63.97
500 62.57 66.20 65.57 62.35 66.94
Table 9: Detailed results of ablation study in the six text classification tasks. We select 100 instances in each iteration and conduct 4 iterations. The results are measured using macro-F1 scores and they are the average values on three different initial sets $D_{l}$ and three different random seeds.
E.4 Capacity of Score
To assess our model’s capability to distinguish between reasonable and unreasonable explanations, we evaluate its ranking performance on the test set. Specifically, after four iterations of the AL process as per section 5.1, we prompt ChatGPT to generate diverse explanations for the test data and score them using Eq. 6. In each test step, we feed both a reasonable and an unreasonable explanation to our model and calculate the accuracy in predicting the reasonable ones based on the computed explanation score (Table 10). As seen in the results, the model incorporating ranking loss achieves superior performance compared to the model without it. For instance, the accuracy in RTE and MRPC are 73.93% and 78.62% in the model with ranking loss, which are 5.36% and 4.30% higher than those without ranking loss, respectively. The improvement in prediction accuracy suggests that the ranking loss can enhance the model’s ability to score the reasonability of explanations.
RTE MRPC MAMS COVID19 DEBA CLEF
w Ranking Loss 73.93 78.62 68.64 62.25 59.78 90.73
w/o Ranking Loss 68.57 74.32 62.04 58.34 55.04 87.39
Table 10: The performance of scoring the explanations w/o ranking loss.
Appendix F Human Annotation
We also carry out experiments to analyze how the human generation of explanations impacts model performance. Specifically, we replace ChatGPT with human annotation to generate explanations on the MAMS dataset. For human annotation, three PhD students specializing in NLP annotate the labels and explanations. Specifically, the models achieve the macro-F1 scores of RTE-62.13%, MRPC-63.36%, MAMS-67.38%, COVID19-69.70%, and CLEF-71.56%, which are relatively lower compared to ChatGPT, which could be due to inconsistent annotation styles among annotators and changes in the annotation scheme from the original dataset (Gilardi et al., 2023; Zhu et al., 2023). The results also demonstrate the effectiveness of explanation generation through LLMs in XAL.
Appendix G Cost Analysis
We conduct experiments (on one GPU V100 Tesla) to analyze the time consumption during each data query process in the MAMS task, which involves 7,090 training data instances. The results are as follows: ME-2 minutes, CA-2 minutes, BK-2 minutes, LC-2 minutes, BALD-11 minutes, Coreset-54 minutes, and our model XAL-21 minutes. Upon observation, it’s apparent that our model requires more time for querying unlabeled data when compared to methods that leverage model uncertainty. However, it consumes less time than the representativeness-based method Coreset.
Appendix H Sensitivity Analysis
In this study, we establish our hyper-parameters based on the relative magnitude and importance of various loss functions, consistently applying these across all datasets without resorting to grid search for optimization. This section further explores the sensitivity of our model to these hyper-parameters. Initially, we examine the impact of different $\lambda_{1}$ values, as depicted in Figure 8. Our observations reveal that the model’s performance remains relatively stable with $\lambda_{1}$ values of 0.3 and 0.1. However, a notable decline in performance occurs when $\lambda_{1}$ is increased to 0.5, attributed to the generative loss becoming approximately ten times greater than the classification loss. Conversely, reducing $\lambda_{1}$ to 0.05 results in a significant deterioration in model performance, suggesting that excessively minimizing the generative loss is detrimental. Subsequently, we assess the model’s response to various $\lambda_{2}$ values, detailed in Figure 9. These findings indicate that higher $\lambda_{2}$ values can adversely affect the model’s performance. Yet, the model exhibits lesser sensitivity to changes in $\lambda_{2}$ when it is equal to or less than 0.01.
Figure 8: The experimental results with different hyper-parameter $\lambda_{1}$ in MAMS.
Figure 9: The experimental results with different hyper-parameter $\lambda_{2}$ in MAMS.
Appendix I Case Study
I.1 Model Generation
Some generation cases are shown (Table 11) from the models trained for 500 data in the AL process of Section 5.1. In these cases, we can find that our model can generate reasonable explanations for the label, which indicates the interpretability of our inference. But in some cases such as the case iv., although the explanation generates the correct label Agree, it explains the label with a wrong reason, which implies that the explainer does not perform perfectly in the small number of data. But it also indicates that we can enhance the model performance in inference and generation by selecting the data with unreasonable explanations through human beings.
Input
Golden
Explanation
i. RTE
Sentence 1: Dana Reeve, the widow of the actor Christopher Reeve, has died of lung cancer at age 44, according to the Christopher Reeve Foundation.
Sentence 2: Christopher Reeve had an accident.
Not Entailment
Sentence 1 states that Dana Reeve, the widow of Christopher Reeve, has died of lung cancer at the age of 44. It does not provide any information about what caused Christopher Reeve’s accident. Therefore, sentence 1 does not entail sentence 2.
ii. MRPC
Sentence 1: Air Commodore Quaife said the Hornets remained on three-minute alert throughout the operation.
Sentence 2: Air Commodore John Quaife said the security operation was unprecedented.
Not Equivalent
The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information. Sentence 1 states that the Hornets were on three-minute alert during the operation, while Sentence 2 states that the security operation was unprecedented.
iii. MAMS
Review: Waiter took our drink order and then we didn’t see him for 15 minutes.
Category: Stuff.
Negative
The reason for the negative sentiment towards the staff in this review is because the waiter took the drink order and disappeared after 15 minutes. This indicates a lack of professionalism and attentiveness, which can be frustrating for customers.
iv. COVID19
Tweet: @WCheesemaker @StephenKing No. You’re missing the point. And from the way I phrased my statement I can understand why. I was attempting to point out masks work! They help stop the spread of disease. I should have stated it more clearly. I’m a scientist; a big fan of masks- and science!! :relaxed:
Target: Face mask.
Favor
The stance of this tweet is in favor of face masks. The reason given is that the tweeter believes masks are effective in stopping the spread of disease, and therefore should be encouraged to use them.
v. DEBA
Sentence 1: Best thing you can do is simply not drive. Take public transportation instead.
Sentence 2: That’s hugely unrealistic for many people living in places with poor public transport provisions (rural areas, poorer areas) or who live in poorly planned towns where amenities are spatially distant.
Disagree
Sentence 2 disagrees with Sentence 1 because it argues that taking public transportation instead of driving is not feasible for many people living in places with poor public transport provisions or in poorly planned towns.
vi. DEBA
Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters. This is… uh, not a guy who’s gonna accomplish much in office.
Sentence 2: Biden has surprised me with some of his action but on the whole it’s still up to activists and local politicians to get stuff done. Shame
Agree
Sentence 2 agrees with Sentence 1 because it acknowledges that Biden has made some progress, but it also emphasizes that it is up to activists and local politicians to bring about meaningful change.
Table 11: Case study of the explanation generation of our model. The model is trained on 500 labeled data instances following the AL process in Section 5.1.
I.2 Unreasonable Generation
We also show some cases that our model believes have high unreasonability in the training set (Table 12). It is noted that in these cases the model generates some unreasonable explanations.
Input
Golden
Explanation
i. MAMS
Review: I like the smaller portion size for dinner.
Category: miscellaneous
Positive
The reason for the neutral sentiment polarity towards the aspect "miscellaneous" in the review is that the reviewer prefers the smaller portion size for dinner. This suggests that the overall dining experience is more enjoyable and satisfying, resulting in a neutral sentiment polarity.
ii. MAMS
Review: After ordering drinks, we both decided on the Paella Vallenciana, brought out on hot plates.
Category: food
Neutral
The positive sentiment polarity of the review towards the aspect of food is due to the description of the Paella Vallenciana being brought out on hot plates. This suggests that the dish was served fresh and hot, enhancing the dining experience and leaving a positive impression.
iii. COVID19
Tweet: @Timewalkproject @BBCNormanS @BBCNews A lot less palaver than being hooked up to a life support machine! #WearAMask
Target: Face Mask
Favor
The tweet expresses a negative stance towards face masks, referring to them as COVID-19. The reason for this stance is not provided within the given word limit.
iv. DEBA
Sentence 1: Energy mostly controls nuclear weapons and nuclear power plants, not fossil fuels. That’s why they put a nuclear physicist (Muniz) in charge of it. What do people think Muniz going to do that will be harmful to the environment? Is this just an indirect anti-nuclear attack? Or just ignorance by insignificant, fringe environmental groups?
Sentence 2: DoE also has the National Renewable Energy Lab, where a lot of the tech development used for renewable energy has happened. The DoE funded the development of fracking. There’s a history of recent appointees using positions there to champion the fossil fuels industry, even if most of what the DoE does is nuclear.
Agree
Sentence 2 disagrees with sentence 1 because it suggests that governments have not used taxes for the intended purposes. This implies that they raid the education budget to raise taxes for education, contradicting the claim in sentence 1 that the purpose of a carbon tax is to.
v. DEBA
Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters. This is… uh, not a guy who’s gonna accomplish much in office
Sentence 2: And that will please you, because you’d rather complain about it and have your cynicism be validated.
Disagree
Sentence 2 agrees with Sentence 1 because it argues that the person who posts here is not trying to debate, but rather to annoy. It suggests that the solution is to simply out-annoy them and ban them.
Table 12: Case study of the unreasonable explanation of our model. The model is trained on 500 labeled data instances following the AL process in Section 5.1.
Appendix J Representation Visualization
We further demonstrate more visualizations (Figure 10) in DEBA and Covid19 to show the effectivness of XAL in exploring informative data.
Figure 10: The t-SNE visualization of sentence representations in the data selection process.

### v3

@@
 Recently, LLMs of generative schema have shown excellent performance in various NLP tasks.
-However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021).
+However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), inclination to their internal knowledge Yan et al.
+(2024), catastrophic forgetting during instruction tuning Luo et al.
+(2023b), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021).
 Considering the problems, LLMs are also applied to improve the generalization of smaller models for specific tasks.
@@
 3.4 Generation of Golden Explanations
-Previous work has shown that LLMs are good at reasoning (Bang et al., 2023; Rajasekharan et al., 2023).
+Previous work has shown that LLMs are good at reasoning (Bang et al., 2023; Rajasekharan et al., 2023; Hsieh et al., 2023).
 Inspired by these studies, we take the LLMs, such as ChatGPT and GPT4, as the teacher models, and query them to generate explanations for each selected labeled data, eliminating the annotation cost of human labor.
@@
 4 Experiments
-Figure 3: Results given the data selection budget 500 instances in six text classification tasks, where 100 instances are selected for annotation in each iteration.
-Here we plot the specific values of XAL and the second significant performance when using 500 instances, and the detailed performance values can be found in Appendix E.1.
+Task
+Dataset
+# Labels Train Dev Test
+Natural Language Inference
+RTE (Bentivogli et al., 2009)
+2 2,240 250 278
+Paraphrase Detection
+MRPC (Dolan et al., 2004)
+2 3,667 409 1,726
+Stance Detection
+COVID19 (Glandt et al., 2021)
+3 4,533 800 800
+Category Sentiment Classification
+MAMS (Jiang et al., 2019)
+3 7,090 888 901
+(Dis)agreement Detection
+DEBA (Pougué-Biyong et al., 2021)
+3 4,617 578 580
+Relevance Classification
+CLEF (Kanoulas et al., 2017)
+2 7,847 981 982
+Table 1: All the six text classification tasks used in our experiments.
+The extent of difficulty is roughly in an increasing tendency.
+RTE and MRPC are fundamental natural language tasks and are included in the widely used benchmark GLUE.
+MAMS, COVID19, and DEBA require the model to understand the text and give suitable inferences towards a specific target or text, and CLEF further provides a difficult dataset with imbalanced label distribution.
 4.1 Tasks and Dataset
-We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text; (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff; (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target; (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment; (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic.
-We adopt the datasets RTE Bentivogli et al.
-(2009), MRPC Dolan et al.
-(2004), MAMS Jiang et al.
-(2019), COVID19 Glandt et al.
-(2021), DEBA Pougué-Biyong et al.
-(2021) and CLEF Kanoulas et al.
-(2017) for each task, respectively.
+We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text Bentivogli et al.
+(2009); Wang et al.
+(2018); (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff Jiang et al.
+(2019); Luo et al.
+(2022a); (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target Glandt et al.
+(2021); Luo et al.
+(2022b); (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment Pougué-Biyong et al.
+(2021); Luo et al.
+(2023a); (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic Kanoulas et al.
+(2017); Guo et al.
+(2023).
 Appendix A demonstrates the details of six datasets with examples.
@@
 (2021) encourages diversity in the vector of expected changes in scores for unlabelled data; (9) Contrastive Active Learning (CAL) (Margatina et al., 2021) selects instances with the maximum mean Kullback-Leibler (KL) divergence between its $m$ nearest neighbors.
+Figure 3: Results given the data selection budget 500 instances in six text classification tasks, where 100 instances are selected for annotation in each iteration.
+Here we plot the specific values of XAL and the second significant performance when using 500 instances, and the detailed performance values can be found in Appendix E.1.
 5 Results and Discussion
@@
 Appendix A Tasks and Corresponding Prompts
-We first introduce the dataset and its statistics in Table 3.
 We show the tasks and examples for experiments in Table 4, including natural language inference, paraphrase detection, category sentiment classification, stance detection, (dis)agreement detection, and relevance classification.
@@
 Figure 7: The process to generate diverse explanations from LLMs. We can obtain reasonable and unreasonable explanations by querying ChatGPT with correct and incorrect labels, respectively.
-Task
-Dataset
-# Labels Train Dev Test
-Natural Language Inference
-RTE (Bentivogli et al., 2009)
-2 2,240 250 278
-Paraphrase Detection
-MRPC (Dolan et al., 2004)
-2 3,667 409 1,726
-Stance Detection
-COVID19 (Glandt et al., 2021)
-3 4,533 800 800
-Category Sentiment Classification
-MAMS (Jiang et al., 2019)
-3 7,090 888 901
-(Dis)agreement Detection
-DEBA (Pougué-Biyong et al., 2021)
-3 4,617 578 580
-Relevance Classification
-CLEF (Kanoulas et al., 2017)
-2 7,847 981 982
-Table 3: All the six text classification tasks used in our experiments.
-The extent of difficulty is roughly in an increasing tendency.
-RTE and MRPC are fundamental natural language tasks and are included in the widely used benchmark GLUE.
-MAMS, COVID19, and DEBA require the model to understand the text and give suitable inferences towards a specific target or text, and CLEF further provides a difficult dataset with imbalanced label distribution.
 Task

### v4

@@
 XAL: EXplainable Active Learning Makes Classifiers Better Low-resource Learners
+Yun Luo1,3    Zhen Yang2    Fandong Meng2    Yingjie Li 1    Fang Guo1,3
 […]
+1School of Engineering, Westlake University   2WeChat AI, Tencent Inc.
+3 Zhejiang University 4 School of Cyber Science and Engineering, Sichuan University
+5 Institute of Advanced Technology, Westlake Institute for Advanced Study
+{luoyun,zhangyue}@westlake.edu.cn
 1 Introduction
@@
 To our knowledge, we are the first to incorporate the model’s explanation (explanation score) to improve the effectiveness of data selection in AL process.
-We will release the code and data to facilitate further research.
+The codes and data have been released in the link to facilitate further research  [^1] https://github.com/LuoXiaoHeics/XAL.
 2 Related Work
@@
 Large Language Model.
-Recently, LLMs of generative schema have shown excellent performance in various NLP tasks.
+Recently, LLMs of generative schema have shown excellent performance in various NLP tasks bubeck2023sparks,yang2023teal.
 However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), inclination to their internal knowledge Yan et al.
 (2024), catastrophic forgetting during instruction tuning Luo et al.
-(2023b), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021).
+(2023), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021).
 Considering the problems, LLMs are also applied to improve the generalization of smaller models for specific tasks.
@@
 4 Experiments
-Task
-Dataset
-# Labels Train Dev Test
-Natural Language Inference
-RTE (Bentivogli et al., 2009)
-2 2,240 250 278
-Paraphrase Detection
-MRPC (Dolan et al., 2004)
-2 3,667 409 1,726
-Stance Detection
-COVID19 (Glandt et al., 2021)
-3 4,533 800 800
-Category Sentiment Classification
-MAMS (Jiang et al., 2019)
-3 7,090 888 901
-(Dis)agreement Detection
-DEBA (Pougué-Biyong et al., 2021)
-3 4,617 578 580
-Relevance Classification
-CLEF (Kanoulas et al., 2017)
-2 7,847 981 982
+Task Dataset # Labels Train Dev Test
+Natural Language Inference RTE (Bentivogli et al., 2009) 2 2,240 250 278
+Paraphrase Detection MRPC (Dolan et al., 2004) 2 3,667 409 1,726
+Stance Detection COVID19 (Glandt et al., 2021) 3 4,533 800 800
+Category Sentiment Classification MAMS (Jiang et al., 2019) 3 7,090 888 901
+(Dis)agreement Detection DEBA (Pougué-Biyong et al., 2021) 3 4,617 578 580
+Relevance Classification CLEF (Kanoulas et al., 2017) 2 7,847 981 982
 Table 1: All the six text classification tasks used in our experiments.
@@
 4.1 Tasks and Dataset
-We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text Bentivogli et al.
-(2009); Wang et al.
-(2018); (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff Jiang et al.
-(2019); Luo et al.
-(2022a); (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target Glandt et al.
-(2021); Luo et al.
-(2022b); (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment Pougué-Biyong et al.
-(2021); Luo et al.
-(2023a); (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic Kanoulas et al.
-(2017); Guo et al.
-(2023).
-Appendix A demonstrates the details of six datasets with examples.
+We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text; (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff; (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target; (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment; (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic.
+The details of the dataset we used are shown in Table 1.
+Appendix A demonstrates the details and prompts of six datasets with examples.
 [^2] Without losing generality, we randomly split the training set in RTE, and MRPC into train/dev set with proportion 9:1.
@@
 We utilize FLAN-T5-large Chung et al.
-(2022) as our backbone network 222 We implement baselines using FLAN-T5-large as well., ChatGPT is adopted to generate the explanations.
+(2022) as our backbone network 333 We implement baselines using FLAN-T5-large as well., ChatGPT is adopted to generate the explanations.
 The implemented details can be found in Appendix D and the detailed values of the result can be found in Appendix E.1.
@@
 The results are presented in Table 2.
-We also observe that the ALPACA-7B can also provide useful explanations to some extent and enhance the model performance compared with ME.
+We also observe that the ALPACA-7B can also provide useful explanations to some extent and enhance the model performance compared with ME through our framework.
 This suggests that LLMs, when used as an assistant in XAL, can provide consistent explanation generation and enhance model performance.
@@
 We honor the ACL Code of Ethics. No private data or non-public information was used in this work. All annotators have received labor fees corresponding to the amount of their annotated instances.
+9 Acknowledgement
+We acknowledge financial support of the National Natural Science Foundation of China Key Program under Grant Number 62336006.
 Appendix A Tasks and Corresponding Prompts
@@
 Figure 7: The process to generate diverse explanations from LLMs. We can obtain reasonable and unreasonable explanations by querying ChatGPT with correct and incorrect labels, respectively.
-Task
-Text
-Natural Language Inference
-Sentence 1: Danny Kennedy, Greenpeace campaigns director, said: "The burden of proof in the Scott Parkin expulsion case lies morally with the Commonwealth, to prove that he is a danger.
+Task Text
+Natural Language Inference Sentence 1: Danny Kennedy, Greenpeace campaigns director, said: "The burden of proof in the Scott Parkin expulsion case lies morally with the Commonwealth, to prove that he is a danger.
 When the Government brought in anti-terror legislation, they promised the public that these laws would only be used to confront a real and present risk of a terrorist attack, not a sweep-all approach against citizens.
@@
 Peace is not a threat to national security.
-No democratic government should expel a foreign citizen because [it] opposes his political opinions."
-Sentence 2: Greenpeace director said that peace is terrorism.
+No democratic government should expel a foreign citizen because [it] opposes his political opinions." Sentence 2: Greenpeace director said that peace is terrorism.
 Label: Not Entailment.
-Paraphrase Detection
-Sentence 1: Last week the power station’s US owners, AES Corp, walked away from the plant after banks and bondholders refused to accept its financial restructuring offer .",
-Sentence 2: "The news comes after Drax’s American owner, AES Corp. AES.N, last week walked away from the plant after banks and bondholders refused to accept its restructuring offer.
+Paraphrase Detection Sentence 1: Last week the power station’s US owners, AES Corp, walked away from the plant after banks and bondholders refused to accept its financial restructuring offer .", Sentence 2: "The news comes after Drax’s American owner, AES Corp.
+AES.N, last week walked away from the plant after banks and bondholders refused to accept its restructuring offer.
 Label: Paraphrase/Semantic Equivalent.
-Category Sentiment Classification
-Text: I left feeling unsatisfied, except for having a nice chance to people watch in the cozy atmosphere with my over-priced pasta bolognese.
-Target: Ambience
-Label: Positive
-Stance Detection
-Text: Michigan is fining individuals 500\$ for not wearing a mask in public.
+Category Sentiment Classification Text: I left feeling unsatisfied, except for having a nice chance to people watch in the cozy atmosphere with my over-priced pasta bolognese.
+Target: Ambience Label: Positive
+Stance Detection Text: Michigan is fining individuals 500\$ for not wearing a mask in public.
 How do y’all feel about this?
@@
 #MaskOn.
-Target: Face Mask
-Label: Favor
-(Dis)agreement Detection
-Text 1: True, but with lower power usage, you have less heat to dissipate, meaning you can overclock it even more.
-Text 2: AMD creates a chip that saves energy by over 31 times. Someone show this to r/PCMasterRace cause we need to switch to AMD.
+Target: Face Mask Label: Favor
+(Dis)agreement Detection Text 1: True, but with lower power usage, you have less heat to dissipate, meaning you can overclock it even more.
+Text 2: AMD creates a chip that saves energy by over 31 times.
+Someone show this to r/PCMasterRace cause we need to switch to AMD.
 Label: Agree.
-Relevance Classification
-Document: 99mtechnetium penicillamine: a renal cortical scanning agent.
+Relevance Classification Document: 99mtechnetium penicillamine: a renal cortical scanning agent.
 99mTechnetium penicillamine, a renal cortical imaging agent, can be used to provide a rapid, safe, and non-invasive assessment of renal morphology and the renal vascular supply.
@@
 Table 4: Tasks and examples for experiments.
-Task
-Prompts
-Label set
-Natural Language Inference
-Sentence 1: {Text 1}. Sentence 2: {Text 2}. Sentence 1 can {Label} sentence 2, explain the reason within 50 words.
-{Entail, Not Entail}
-Paraphrase Detection
-Sentence 1: {Text 1}. Sentence 2: {Text 2}. The relation between the above two sentences is {Label}, explain the reason within 50 words.
+Task Prompts Label set
+Natural Language Inference Sentence 1: {Text 1}. Sentence 2: {Text 2}. Sentence 1 can {Label} sentence 2, explain the reason within 50 words. {Entail, Not Entail}
+Paraphrase Detection Sentence 1: {Text 1}.
+Sentence 2: {Text 2}.
+The relation between the above two sentences is {Label}, explain the reason within 50 words.
 {Paraphrase Equivalent, Not Paraphrase Equivalent}
-Category Sentiment Classification
-{Text}.
-{Positive, Negative, Neutral}
-Stance Detection
-{Text}. The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words.
-{Favor, Neutral, Against}
-(Dis)agreement Detection
-Sentence 1: {Text 1}. Sentence 2: {Text 2}. The stance of sentence 2 is {Label} to sentence 1, explain the reason within 50 words.
-{Agree, Neutral, Disagree}
-Relevance Classification
-{Document}. The scientific document is {Label}’ to the research topic {Topic}, explain the reason within 100 words.
-{Relevant, Not Relevant}
+Category Sentiment Classification {Text}. {Positive, Negative, Neutral}
+Stance Detection {Text}. The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words. {Favor, Neutral, Against}
+(Dis)agreement Detection Sentence 1: {Text 1}. Sentence 2: {Text 2}. The stance of sentence 2 is {Label} to sentence 1, explain the reason within 50 words. {Agree, Neutral, Disagree}
+Relevance Classification {Document}. The scientific document is {Label}’ to the research topic {Topic}, explain the reason within 100 words. {Relevant, Not Relevant}
 Table 5: The prompts we adopted for the tasks in our experiments.
@@
 Using the prompts in Appendix A, we show some examples of the obtained explanations in Table 6 by querying ChatGPT.
-Input
-Label
-Explanation
-RTE
-Sentence 1: The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes.
+Input Label Explanation
+RTE Sentence 1: The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes.
 Sentence 2: Nobel Peace Prize candidates have been chosen.
-Not Entailment
-Sentence 1 states that the Nobel prizes are the main forms of recognition for various fields, including peace.
+Not Entailment Sentence 1 states that the Nobel prizes are the main forms of recognition for various fields, including peace.
 However, sentence 2 only mentions that candidates have been chosen for the Nobel Peace Prize, without specifying if they have actually received the prize.
 Therefore, sentence 2 does not necessarily follow from sentence 1.
-Entailment
-Sentence 1 can entail sentence 2 because sentence 1 mentions that one of the main forms of recognition for those working for peace is the Nobel prizes.
+Entailment Sentence 1 can entail sentence 2 because sentence 1 mentions that one of the main forms of recognition for those working for peace is the Nobel prizes.
 Therefore, it is reasonable to assume that candidates for the Nobel Peace Prize have been chosen.
-MRPC
-Sentence 1: Maddox, 87, cracked two ribs when he fell about 10 days ago at an assisted living home where he was recovering from intestinal surgery, Virginia Carnes said.
+MRPC Sentence 1: Maddox, 87, cracked two ribs when he fell about 10 days ago at an assisted living home where he was recovering from intestinal surgery, Virginia Carnes said.
 Sentence 2: Maddox, who had battled cancer since 1983, cracked two ribs earlier this month when he fell at an assisted living home where he was recovering from surgery.
-Equivalent
-The two sentences provide similar information but are expressed differently.
+Equivalent The two sentences provide similar information but are expressed differently.
 Sentence 1 mentions that Maddox fell about 10 days ago, while Sentence 2 specifies that the fall occurred earlier this month.
 Despite the slight difference in wording, both sentences convey the same meaning.
-Not Equivalent
-The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information about the timing and reason for Maddox’s rib injury.
+Not Equivalent The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information about the timing and reason for Maddox’s rib injury.
 In the first sentence, it mentions that Maddox fell about 10 days ago while recovering from intestinal surgery, whereas the second sentence states that he fell earlier this month while recovering from surgery without specifying the exact time frame.
-COVID19
-Tweet: Placed2nd @AngelaBelcamino Omg people say I have to wear a mask and I’m crying because I have to wear a mask to save lives boo hoo!! MAGAts are stupid!!!
-Favor
-The stance of this tweet towards face masks is favorable.
+COVID19 Tweet: Placed2nd @AngelaBelcamino Omg people say I have to wear a mask and I’m crying because I have to wear a mask to save lives boo hoo!!
+MAGAts are stupid!!!
+Favor The stance of this tweet towards face masks is favorable.
 The reason for this is that the tweet expresses support for wearing masks to save lives and criticizes those who oppose wearing masks, referring to them as ""MAGAts"" (a term often used to describe supporters of former President Donald Trump)."
-Against
-The reason behind the stance against face masks in this tweet seems to be frustration and ridicule towards those who oppose wearing masks.
+Against The reason behind the stance against face masks in this tweet seems to be frustration and ridicule towards those who oppose wearing masks.
 The tweet suggests that the person finds the idea of wearing masks to save lives unnecessary and mocks those who support it, particularly referring to them as ""MAGAts,"" a derogatory term for supporters of former President Trump.
-Neutral
-The tweet expresses a neutral stance towards face masks by mocking the idea of crying over having to wear one.
+Neutral The tweet expresses a neutral stance towards face masks by mocking the idea of crying over having to wear one.
 While it criticizes MAGA supporters, it does not explicitly state a negative or positive opinion about face masks themselves."
@@
 But it also indicates that we can enhance the model performance in inference and generation by selecting the data with unreasonable explanations through human beings.
-Input
-Golden
-Explanation
-i. RTE
-Sentence 1: Dana Reeve, the widow of the actor Christopher Reeve, has died of lung cancer at age 44, according to the Christopher Reeve Foundation.
+Input Golden Explanation
+i.
+RTE Sentence 1: Dana Reeve, the widow of the actor Christopher Reeve, has died of lung cancer at age 44, according to the Christopher Reeve Foundation.
 Sentence 2: Christopher Reeve had an accident.
-Not Entailment
-Sentence 1 states that Dana Reeve, the widow of Christopher Reeve, has died of lung cancer at the age of 44.
+Not Entailment Sentence 1 states that Dana Reeve, the widow of Christopher Reeve, has died of lung cancer at the age of 44.
 It does not provide any information about what caused Christopher Reeve’s accident.
 Therefore, sentence 1 does not entail sentence 2.
-ii. MRPC
-Sentence 1: Air Commodore Quaife said the Hornets remained on three-minute alert throughout the operation.
+ii.
+MRPC Sentence 1: Air Commodore Quaife said the Hornets remained on three-minute alert throughout the operation.
 Sentence 2: Air Commodore John Quaife said the security operation was unprecedented.
-Not Equivalent
-The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information.
+Not Equivalent The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information.
 Sentence 1 states that the Hornets were on three-minute alert during the operation, while Sentence 2 states that the security operation was unprecedented.
-iii. MAMS
-Review: Waiter took our drink order and then we didn’t see him for 15 minutes.
+iii.
+MAMS Review: Waiter took our drink order and then we didn’t see him for 15 minutes.
 Category: Stuff.
-Negative
-The reason for the negative sentiment towards the staff in this review is because the waiter took the drink order and disappeared after 15 minutes.
+Negative The reason for the negative sentiment towards the staff in this review is because the waiter took the drink order and disappeared after 15 minutes.
 This indicates a lack of professionalism and attentiveness, which can be frustrating for customers.
-iv. COVID19
-Tweet: @WCheesemaker @StephenKing No.
+iv.
+COVID19 Tweet: @WCheesemaker @StephenKing No.
 You’re missing the point.
@@
 I’m a scientist; a big fan of masks- and science!!
-:relaxed:
-Target: Face mask.
-Favor
-The stance of this tweet is in favor of face masks. The reason given is that the tweeter believes masks are effective in stopping the spread of disease, and therefore should be encouraged to use them.
-v. DEBA
-Sentence 1: Best thing you can do is simply not drive. Take public transportation instead.
+:relaxed: Target: Face mask.
+Favor The stance of this tweet is in favor of face masks.
+The reason given is that the tweeter believes masks are effective in stopping the spread of disease, and therefore should be encouraged to use them.
+v.
+DEBA Sentence 1: Best thing you can do is simply not drive.
+Take public transportation instead.
 Sentence 2: That’s hugely unrealistic for many people living in places with poor public transport provisions (rural areas, poorer areas) or who live in poorly planned towns where amenities are spatially distant.
-Disagree
-Sentence 2 disagrees with Sentence 1 because it argues that taking public transportation instead of driving is not feasible for many people living in places with poor public transport provisions or in poorly planned towns.
-vi. DEBA
-Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters.
+Disagree Sentence 2 disagrees with Sentence 1 because it argues that taking public transportation instead of driving is not feasible for many people living in places with poor public transport provisions or in poorly planned towns.
+vi.
+DEBA Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters.
 This is… uh, not a guy who’s gonna accomplish much in office.
-Sentence 2: Biden has surprised me with some of his action but on the whole it’s still up to activists and local politicians to get stuff done. Shame
-Agree
-Sentence 2 agrees with Sentence 1 because it acknowledges that Biden has made some progress, but it also emphasizes that it is up to activists and local politicians to bring about meaningful change.
+Sentence 2: Biden has surprised me with some of his action but on the whole it’s still up to activists and local politicians to get stuff done.
+Shame Agree Sentence 2 agrees with Sentence 1 because it acknowledges that Biden has made some progress, but it also emphasizes that it is up to activists and local politicians to bring about meaningful change.
 Table 11: Case study of the explanation generation of our model. The model is trained on 500 labeled data instances following the AL process in Section 5.1.
@@
 We also show some cases that our model believes have high unreasonability in the training set (Table 12). It is noted that in these cases the model generates some unreasonable explanations.
-Input
-Golden
-Explanation
-i. MAMS
-Review: I like the smaller portion size for dinner.
-Category: miscellaneous
-Positive
-The reason for the neutral sentiment polarity towards the aspect "miscellaneous" in the review is that the reviewer prefers the smaller portion size for dinner.
+Input Golden Explanation
+i.
+MAMS Review: I like the smaller portion size for dinner.
+Category: miscellaneous Positive The reason for the neutral sentiment polarity towards the aspect "miscellaneous" in the review is that the reviewer prefers the smaller portion size for dinner.
 This suggests that the overall dining experience is more enjoyable and satisfying, resulting in a neutral sentiment polarity.
-ii. MAMS
-Review: After ordering drinks, we both decided on the Paella Vallenciana, brought out on hot plates.
-Category: food
-Neutral
-The positive sentiment polarity of the review towards the aspect of food is due to the description of the Paella Vallenciana being brought out on hot plates.
+ii.
+MAMS Review: After ordering drinks, we both decided on the Paella Vallenciana, brought out on hot plates.
+Category: food Neutral The positive sentiment polarity of the review towards the aspect of food is due to the description of the Paella Vallenciana being brought out on hot plates.
 This suggests that the dish was served fresh and hot, enhancing the dining experience and leaving a positive impression.
-iii. COVID19
-Tweet: @Timewalkproject @BBCNormanS @BBCNews A lot less palaver than being hooked up to a life support machine! #WearAMask
-Target: Face Mask
-Favor
-The tweet expresses a negative stance towards face masks, referring to them as COVID-19. The reason for this stance is not provided within the given word limit.
-iv. DEBA
-Sentence 1: Energy mostly controls nuclear weapons and nuclear power plants, not fossil fuels.
+iii.
+COVID19 Tweet: @Timewalkproject @BBCNormanS @BBCNews A lot less palaver than being hooked up to a life support machine!
+#WearAMask Target: Face Mask Favor The tweet expresses a negative stance towards face masks, referring to them as COVID-19.
+The reason for this stance is not provided within the given word limit.
+iv.
+DEBA Sentence 1: Energy mostly controls nuclear weapons and nuclear power plants, not fossil fuels.
 That’s why they put a nuclear physicist (Muniz) in charge of it.
@@
 There’s a history of recent appointees using positions there to champion the fossil fuels industry, even if most of what the DoE does is nuclear.
-Agree
-Sentence 2 disagrees with sentence 1 because it suggests that governments have not used taxes for the intended purposes.
+Agree Sentence 2 disagrees with sentence 1 because it suggests that governments have not used taxes for the intended purposes.
 This implies that they raid the education budget to raise taxes for education, contradicting the claim in sentence 1 that the purpose of a carbon tax is to.
-v. DEBA
-Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters.
-This is… uh, not a guy who’s gonna accomplish much in office
-Sentence 2: And that will please you, because you’d rather complain about it and have your cynicism be validated.
-Disagree
-Sentence 2 agrees with Sentence 1 because it argues that the person who posts here is not trying to debate, but rather to annoy. It suggests that the solution is to simply out-annoy them and ban them.
+v.
+DEBA Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters.
+This is… uh, not a guy who’s gonna accomplish much in office Sentence 2: And that will please you, because you’d rather complain about it and have your cynicism be validated.
+Disagree Sentence 2 agrees with Sentence 1 because it argues that the person who posts here is not trying to debate, but rather to annoy.
+It suggests that the solution is to simply out-annoy them and ban them.
 Table 12: Case study of the unreasonable explanation of our model. The model is trained on 500 labeled data instances following the AL process in Section 5.1.